tom is twice as old as harry in a years time their combine ages will be 5 time. harry,present age how old is tomorrow ? how old will be in 9 years time ?
Answer by stanbon(75887)

tom is twice as old as harry.
t = 2h
-----------------
in a years time their combine ages will be 5 time harry,present age.
(t+1)+(h+1) = 5h
-------------------
how old is tomorrow ?
Substitute 2h for t in the 2nd equation to solve for "h":
(2h+1)+(h+1)= 5h
3h + 2 = 5h
2h = 2
h = 1 (Harry's age now)
------------
Since t=2h, t=2 (Tom's age now)
----------------------------------
how old will be in 9 years time ?
Harry will be 1+9 or 10 yrs. old
Tome will be 2+9 or 11 yrs. old
========================================
Cheers,
Stan H.